### v1

# Chen Xu1,2, Zhenyu Lv1,2, Tian Lan4, Xianyang Wang1,2, Luyao Ji3, Leyang Cui Minqiang Yang5, Jian Shen1,2, Qunxi Dong1,2, Xiuling Liu6, Juan Wang3, Bin Hu1,2,5

Abstract
Although large language models (LLMs) hold significant promise in psychotherapy, their direct application in patient-facing scenarios raises ethical and safety concerns. Therefore, this work shifts towards developing an LLM as a supervisor to train real therapists. In addition to the privacy of clinical therapist training data, a fundamental contradiction complicates the training of therapeutic behaviors: clear feedback standards are necessary to ensure a controlled training system, yet there is no absolute “gold standard” for appropriate therapeutic behaviors in practice. In contrast, many common therapeutic mistakes are universal and identifiable, making them effective triggers for targeted feedback that can serve as clearer evidence. Motivated by this, we create a novel therapist-training paradigm: (1) guidelines for mistaken behaviors and targeted correction strategies are first established as standards; (2) a human-in-the-loop dialogue-feedback dataset is then constructed, where a mistake-prone agent intentionally makes standard mistakes during interviews naturally, and a supervisor agent locates and identifies mistakes and provides targeted feedback; (3) after fine-tuning on this dataset, the final supervisor model is provided for real therapist training. The detailed experimental results of automated, human and downstream assessments demonstrate that models fine-tuned on our dataset Mate, can provide high-quality feedback according to the clinical guideline, showing significant potential for the therapist training scenario.
Introduction
Figure 1: A real example from the LLM Supervisor trained on the proposed Mate dataset, where the model draws on a guideline (mistake-correction pairs shown in Figure 2a) as evidence to highlight mistaken utterances, categorize mistaken behaviors, and provide targeted corrective feedback.
Background
The field of mental health is grappling with a profound imbalance between the increasing number of patients and the scarcity of available therapists (WHO 2023). Although Large language models (LLMs) offer a potential solution to this imbalance by simulating human therapist behaviors and providing personalized counseling services at lower costs (Xiao et al. 2024; Na 2024; Xie et al. 2024; Wang et al. 2025) , their direct application in patient-facing treatment poses several risks. These risks include ethical dilemmas (Ong et al. 2024), safety concerns (Hager et al. 2024), potential harm with uncertainty about responsibility (Nichol et al. 2023), and the inherent challenge of establishing a therapeutic alliance with virtual therapists (Choudhury and Chaudhry 2024). This prompts us to explore automated methods for training or supervising therapists to accelerate their development, addressing the imbalance by empowering more therapists through AI-assisted means.
AI-Partner Training Paradigm
With significant advances in LLM role play  (Xu et al. 2022; Tseng et al. 2024; Chen et al. 2024), a new AI-assisted training paradigm has emerged (Yang et al. 2024). In this paradigm, LLMs play partners in the field and collaborate with practitioners to create interactive scenarios for experiential learning (Kolb 2014). For example, in educational scenarios, “AI Student as Partner” helps teaching assistants improve their communication and classroom management skills (Markel et al. 2023); in conflict resolution, “AI Opponent as Partner” helps users practice negotiation and conflict resolution (Shaikh et al. 2023). However, in psychotherapy, “AI Patient as Partner” (Zhang et al. 2025) is difficult to implement. Privacy concerns and limited clinical data hinder the accurate modeling of patients’ behavioral patterns, as well as the already intricate cognitive states (Reger et al. 2021; Lee et al. 2024a).
AI-Patient as Partner
Recently, a pioneering work, Patient-$\Psi$-Trainer has been proposed to fill this gap (Wang et al. 2024). It integrates 106 cognitive models of patients (Beck 2020), manually created by clinical experts, into LLMs to simulate how real patients react through the interplay of beliefs, thoughts, emotions, situations and other factors shown in Figure 2c. With high-fidelity simulations of patients, the ability to understand patients is trained: after interviewing simulated patients, novice therapists are required to independently describe the patient’s underlying cognitive model and receive feedback by comparing it to the standard answer—the cognitive model originally used to simulate their interviewee. This innovative yet safer training paradigm inspires us to explore a more sophisticated and practice-grounded training task aimed at improving the interview behaviors of therapists, specifically their ability to demonstrate effective, adaptive, and appropriate interactive behaviors during interviews, rather than focusing solely on post-session understanding of patients.
Motivation and Main Idea
However, a fundamental contradiction makes the training of therapeutic behaviors particularly challenging: as a training system, more “clear answers” as feedback are required to ensure consistent and goal-oriented learning within a low-risk and controlled environment (Arango-Ibanez et al. 2024)—much like the use of gold-standard cognitive models as feedback in the patient understanding training task. Yet, there is no absolute “gold standard” for appropriate interviewing behaviors in actual therapeutic practice (Leichsenring and Steinert 2017). The diversity and complexity of therapeutic approaches make it very difficult to provide unified, authoritative standards. Imposing a universal standard could lead to homogenized styles, suppress the individuality and creativity of therapists, and undermine their self-efficacy (Larson and Daniels 1998; Leichsenring and Steinert 2017). Conversely, certain mistaken therapy behaviors tend to be universal and relatively easy to identify (Gerke et al. 2020) such as neglect of the patient’s emotional expression, breaches of ethical standards, and transference or countertransference issues (Vyb´ıral et al. 2024). If we can leverage these relatively clear mistaken behaviors as labels to construct the training environment, then when trainees exhibit these behaviors, the corresponding targeted corrective principles can be retrieved and combined with the current context as relatively definite feedback to help them improve—our main idea to move a small step toward such open-ended training tasks.
Proposed Method
Motivated by the above, we present an approach to training LLMs as supervisors that adhere to given standards, aiming to train therapists in a controlled and evidence-based manner. To achieve this, we first collaborated with clinical therapists and supervisors to construct a guideline of 15 common mistakes paired with targeted correction strategies. Building on this guideline as a “coordination protocol”, we propose the dataset Mate (MistAken behaviors to TargEted feedback) through a multi-agent framework, where a mistake-prone therapist agent, a mistake-sensitive client agent, and a mistake-corrective supervisor collaboratively generate dialogue-feedback data, which is reviewed by humans in the loop. Finally, the Supervision model fine-tuned on Mate, illustrated in Figure 1, can pinpoint the mistake location, identify its category, and generate targeted and constructive feedback. Experimental validation demonstrates the effectiveness of this approach across multiple dimensions. We conduct comprehensive evaluations including mistake identification, dialogue feedback generation, professional quality assessment, and downstream tasks. Results show that fine-tuning with our Mate dataset significantly enhances model performance, with notable improvements in core supervisory competencies.
Contributions
(1) We propose the LLM-as-a-Supervisor task, where LLMs locate problematic utterance, classify mistake types, and generate corrective feedback based on the training guideline of mistaken behaviors & correction strategy in the safe and controlled environment. (2) We develop a mistake-driven multi-agent human-in-the-loop data synthesis pipeline that generates realistic therapist-client dialogues with embedded mistakes and corresponding supervisory feedback, resulting in Mate, a high-quality psychological dialogue-feedback dataset for Supervised Fine-Tuning. (3) We demonstrate that LLMs finetuned with Mate dataset significantly enhances domain-specific supervisory capabilities, with our 8B and 14B parameter models outperforming multiple closed-source LLMs on the proposed tasks, effectively transforming generic text generation models into specialized professional feedback tools for therapist education.
Related Works
LLM-as-a-Therapist
Existing work has begun to explore pipelines for creating realistic counselor-client dialogues for realistic counselor simulation. For example, the ESConv (Liu et al. 2021) dataset provides multi-turn emotional support dialogues to train supporter models using various strategies, while Chaszczewicz et al. (2024) augments a subset of ESConv with multi-layer feedback, linking strategy execution to evaluative feedback. Qiu and Lan (2024) employ role-play approaches to train LLMs for emotional support. Other efforts in more clinical scenarios, such as Li et al. (2024) and Shen et al. (2024) employ self-play or retrieval-based strategy to train LLM for the treatment of mental health disorders. However, these foundational client-counselor datasets are designed primarily to simulate supporters in realistic settings, rather than originally intended for training settings. As a result, the systematic customization and incorporation of training objectives and standards are often absent from the initial stages of dataset design.
LLM-as-a-Patient
To train therapists in patient cognitive conceptualization, Patient-$\Psi$-Trainer (Wang et al. 2024) constructed patient cognitive models based on CBT principles, which serve as feedback standards for subsequent training. In this framework, LLMs simulate therapy patients, and after each interview, novice therapists are required to write out the patient’s underlying cognitive model, which is then compared to the standard answer for feedback. We adopt this innovative yet safer training paradigm; however, we explore a therapist training task with less standard answers, focusing on therapists’ generative abilities during the interview rather than merely assessing their understanding after the session.
Methodology
Figure 2: Overview of our human-in-the-loop dataset construction pipeline. A multi-agent framework (e) collaboratively generates dialogue-feedback data through: (b) a therapist agent that naturally integrates predefined therapeutic mistakes into conversations, (c) a patient agent that authentically reacts based on its underlying cognitive model, and (d) a supervisor agent that observes therapist-patient interactions to provide targeted feedback. Clinical therapists and supervisors participate in both (a) guideline creation to ensure therapist training objectives and (f) instance review to ensure overall dataset quality.
Task Definition
We aims to develop a LLM-based psychological counseling supervision system that observes therapist-client dialogues and provides professional feedback for novice therapists. Let $\mathcal{H}=\{u_{1},u_{2},...,u_{n}\}$ denote the dialogue history, where $u_{i}$ represents the utterance in the $i$-th turn. Let $\mathcal{S}$ denote the set of mistaken utterances, $\mathcal{M}$ the set of mistake categories, and $\mathcal{F}$ the set of feedback. We formalize the supervision task as the following joint probability distribution:
$P(f,m,s|\mathcal{H})$ (1)
where $f\in\mathcal{F}$, $m\in\mathcal{M}$, and $s\in\mathcal{S}$. According to the chain rule of probability, this joint distribution can be decomposed into three sequential conditional probabilities. First, the model performs Problematic Sentence Location by computing $P(s|\mathcal{H})$, which locates the specific utterance containing the therapist’s mistake based on the dialogue history. Second, given the identified mistaken utterance, the model conducts Mistake Category Classification through $P(m|\mathcal{H},s)$, determining the clinical category of the identified problem. Finally, the model generates comprehensive supervision feedback via Feedback Generation, modeled as $P(f|\mathcal{H},s,m)$, which synthesizes all available information to produce actionable guidance. Therefore, the complete generation process can be expressed as:
$P(f,m,s|\mathcal{H})=P(s|\mathcal{H})\cdot P(m|\mathcal{H},s)\cdot P(f|\mathcal{H},s,m)$ (2)
This formulation enables our system to mimic the cognitive process of real clinical supervisors: observing dialogues, identifying problems, analyzing their nature, and ultimately providing constructive guidance.
Guideline Construction: Mistaken Behaviors & Correction Strategy
To model common therapeutic mistakes of novice therapists, we collaborate with clinical experts to construct Mistaken Behaviors & Correction Strategy Guideline $\mathbb{M}=\{M_{1},M_{2},...,M_{16}\}$, where each element $M_{k}$ is defined as a triple (Figure 2a): $M_{k}=(cat_{k},beh_{k},crit_{k})$ where $cat_{k}$ denotes the category name, $beh_{k}$ provides a detailed description of the mistaken behavior, and $crit_{k}$ specifies the corresponding supervision correction strategy. This set encompasses 15 types of typical mistaken behaviors plus one “Exemplary Practice” category, totaling 16 behavior patterns. This prevents models from becoming overly critical by incorporating positive samples alongside negative ones.
Mistake-Driven Multi-Agent Dialogue-Feedback Generation
Given the scarcity of high-quality supervision data in the psychological counseling domain and the critical challenges of privacy restrictions and sample imbalance in real clinical data, we propose a mistake-driven multi-agent data synthesis pipeline to generate controlled, systematic, and large-scale dialogue-feedback pairs containing specific clinical mistakes (Figure 2e).
Agent Behavior Modeling
The framework is driven by two independent sets: the Mistaken Behavior Guideline $\mathbb{M}$ and the Client Case Set (Wang et al. 2024) $\mathbb{C}=\{C_{1},C_{2},...,C_{106}\}$. For the mistaken behavior therapist $T_{k}$, we instantiate a specialized therapist agent powered by carefully designed prompts for each $M_{k}\in\mathbb{M}$ (Figure 2b). This agent is configured to consistently exhibit the $k$-th specific mistake during conversations. Its behavior in the $i$-th dialogue turn is modeled as:
$u_{i}^{T}=T_{k}(\mathcal{H}_{i-1},cat_{k},beh_{k})$ (3)
where $u_{i}^{T}$ is the therapist’s utterance in the $i$-th turn, $\mathcal{H}_{i-1}=\{u_{1},...,u_{i-1}\}$ represents the dialogue history up to turn i1i-1, and $cat_{k}$ and $beh_{k}$ extracted from $M_{k}$ determine the agent’s behavior pattern. Correspondingly, for each client case $C_{j}\in\mathbb{C}$, we instantiate a behavior-sensitive client agent $C_{j}$ (Figure 2c). This agent possesses unique background and personality traits while being capable of producing contextually appropriate responses to the therapist’s specific mistakes. Its behavior is modeled as:
$u_{i}^{C}=C_{j}(\mathcal{H}_{i-1}\oplus u_{i}^{T},Case_{j})$ (4)
where $u_{i}^{C}$ is the client’s response, $\oplus$ denotes sequence concatenation, $Case_{j}$ contains the client’s information.
Dialogue Generation and Feedback Synthesis
In the dialogue generation stage, the system orchestrates multi-turn interactions between a designated therapist $T_{k}$ and client $C_{j}$, producing a complete dialogue history:
$D_{j,k}=G_{dialogue}(T_{k},C_{j})=\{(u_{1}^{T},u_{1}^{C}),...,(u_{n}^{T},u_{n}^{C})\}$
Following dialogue generation, a clinical supervisor agent $\mathcal{S}$ with an “omniscient perspective” reviews the entire dialogue process $D_{j,k}$ and generates structured critical feedback based on predefined mistake definition $M_{k}$ (Figure 2d):
$F_{j,k}=\mathcal{S}(D_{j,k},M_{k})=(s_{j,k},m_{j,k},f_{j,k})$ (6)
where the feedback $F_{j,k}$ comprises three components: $s_{j,k}$ identifies the problematic sentence by precisely quoting the specific utterance embodying the mistake; $m_{j,k}$ denotes the mistaken category, explicitly stating the category of error committed by the novice counselor; and $f_{j,k}$ provides concrete and actionable feedback such as response strategies and improvement suggestions. Finally, the dialogue and feedback are combined into a complete data pair:
$P_{j,k}=(D_{j,k},F_{j,k})$ (7)
which collectively forms the high-quality dataset required for training our model.
Dataset Construction Details
We employ a full cross-combination strategy, pairing 106 client types with 16 threapy behavior patterns, yielding $|\mathbb{C}|\times|\mathbb{M}|=106\times 16=1,696$ unique combinations. With 6 independent samplings per combination using Deepseek-V3, we generate a comprehensive dataset where each sample is automatically annotated with both mistake category labels (indicating which of the 16 patterns is present) and problematic utterance labels (identifying the specific utterances containing errors).
Data Quality Assurance
To ensure the highest data quality standards, this study implements a comprehensive quality assurance pipeline that processes the initial dataset of 10,176 records through multiple refinement stages (Figure 2f). Through systematic data cleaning, Validator-Guided Refinement (VGR), and expert manual review, the final curated dataset contains 9,915 high-quality data samples.
Validator-Guided Refinement
To ensure data quality, this study introduces a Validator-Guided Refinement (VGR) mechanism that leverages GPT-4o as both validator and refiner to iteratively review and enhance feedback initially generated by the base models through a closed-loop process. The VGR strategy simulates a human expert supervision loop, where the validator model guides the refinement of initial outputs, effectively mitigating issues like hallucination and significantly improving the standardization of the content. Algorithm 1 details the complete process.
Refine Stage: Feedback Optimization. The Refine stage takes the dialogue history $D_{j,k}$ and the initial feedback $F_{j,k}=(s_{j,k},m_{j,k},f_{j,k})$ as input. Through a meticulously designed prompt, it drives the validator model $M_{V}$ to enhance the original feedback $f_{j,k}$, outputting an optimized version $f^{\prime}_{j,k}$.
Validate Stage: Compliance Verification. The refined feedback $f^{\prime}_{j,k}$ generated in the Refine stage is immediately audited by the same validator model $M_{V}$. Employing an “LLM-as-a-judge” framework (Lan et al. 2024), we establish validation criteria across four dimensions(progressivity, actionability, ethicality, supportiveness). The validator takes the dialogue $D_{j,k}$, the original feedback $f_{j,k}$, and the refined feedback $f^{\prime}_{j,k}$ as input, outputting a compliance judgment. The output is considered compliant only when all checklist criteria are met. Otherwise, the algorithm will re-attempt the refinement, up to a maximum of $N_{\text{retry}}$ times. If all attempts fail, the sample is flagged for subsequent human review.
Clinical Expert Manual Refinement
Samples failing automated VGR undergo manual refinement by two licensed clinical psychology experts following a structured protocol: (1) independent review of dialogue and feedback; (2) identification of clinical inaccuracies or logical inconsistencies; (3) collaborative generation of refined feedback through expert consensus. For quality assurance, we randomly sample 5% of the data for expert validation. Inter-rater agreement between automated and manual refinement achieves Cohen’s kappa $\kappa=0.87$, indicating substantial agreement and validating VGR mechanism’s reliability.
Algorithm 1 The Validator-Guided Refinement (VGR) strategy for feedback refinement and validation.
1:Generator Model $M_{G}$, Validator Model $M_{V}$
2:Initial dataset $\mathcal{D}=\{P_{j,k}\}_{j,k=1}^{N}$, where each $P_{j,k}=(D_{j,k},F_{j,k})$
3:Dataset $\mathcal{D}_{\text{refined}}$ with refined feedback components $f^{\prime}_{j,k}$
4:Hyperparameters: $N_{\text{retry}}\in\mathbb{N},T\in(0,1]$
[…]
6:for each $P_{j,k}\in\mathcal{D}$ do
7:  $\textit{refine\_prompt}\leftarrow\textit{BuildPrompt}(D_{j,k},F_{j,k})$
8:  $\textit{flag}\leftarrow\textit{false}$
9:  for $r=1,2,\dots,N_{\text{retry}}$ do
10:   $f^{\prime}_{j,k}\leftarrow\textit{CallLLM}(M_{V},\textit{refine\_prompt},T)$
11:   $\textit{val\_result}\leftarrow\textit{Validator}(M_{V},D_{j,k},f_{j,k},f^{\prime}_{j,k})$
12:   if val_result passes all criteria then
13:     $F^{\prime}_{j,k}\leftarrow(s_{j,k},m_{j,k},f^{\prime}_{j,k})$; $P^{\prime}_{j,k}\leftarrow(D_{j,k},F^{\prime}_{j,k})$
14:     $\textit{flag}\leftarrow\textit{true}$; break
15:   end if
16:  end for
17:  if $\textit{flag}=\textit{false}$ then
18:   $P^{\prime}_{j,k}\leftarrow(D_{j,k},(s_{j,k},m_{j,k},\bot))$ with flag “need_human”
19:  end if
20:  $\mathcal{D}_{\text{refined}}\leftarrow\mathcal{D}_{\text{refined}}\cup\{P^{\prime}_{j,k}\}$
21:end for
22:return $\mathcal{D}_{\text{refined}}$
Dataset Statistics
The final curated dataset comprises 9,915 total records, which are randomly split into training (90%) and test sets (10%). The refinement process statistics reveal the effectiveness of the quality assurance pipeline: 9,386 samples of the original data required no refinement, indicating high initial quality from the base generator models; 5.3% (529 samples) underwent successful automated refinement via VGR; and 1.1% (111 samples) required manual expert intervention. Additionally, 261 samples were filtered out due to fundamental quality issues. This distribution demonstrates the efficiency of our multi-tiered approach, where automated methods handle the majority of cases while preserving human expertise for the most challenging scenarios. Table 1 presents comprehensive statistics.
Metric Train Set Test Set
Dataset Size
Total Records 8,923 992
Diversity Stats
Unique Patient IDs 106 106
Behavior Categories 16 16
Refinement Status
No Refinement Needed 8,441 945
VGR Refined 482 47
Manual Expert Review 102 9
Content Statistics
Avg. Feedback Chars 1,969.85 1,960.14
Avg. Dialogue Turns 12.01 12.08
Avg. Utterance Length 158.15 158.95
Avg. Labeled Prob. Utt. 2.71 2.74
Table 1: Key Statistics of the Mate.
Model Task 1: Mistake Category Classification Task 2: Problematic Sentence Location
Accuracy W. Precision W. Recall W. F1-Score M. Precision M. Recall M. F1-Score M. Jaccard EMR
Closed-Source Models
GPT-4o 51.18 62.69 51.18 45.66 62.19 89.93 69.48 57.63 13.68
Claude-Sonnet-4 54.59 65.82 54.59 47.44 59.57 94.67 70.04 58.04 13.14
Open-Source Models
DeepSeek-V3-0324-685B 52.14 64.04 52.14 46.85 60.71 97.05 71.85 59.66 13.57
Llama3.1-8B 29.39 43.73 29.39 25.96 59.40 94.31 69.91 57.28 10.01
Llama3.1-8B (w/ Mate) 43.81 54.51 43.81 39.64 59.57 94.37 70.27 58.12 13.13
Llama3.1-8B ($\Delta$) +14.42 +10.78 +14.42 +13.68 +0.17 +0.06 +0.36 +0.84 +3.12
[1pt/2pt] Qwen3-8B 23.33 44.16 23.33 22.18 54.78 60.76 53.56 40.52 3.87
Qwen3-8B (w/ Mate) 27.03 49.35 27.03 27.72 58.40 84.53 64.59 51.79 8.12
Qwen3-8B ($\Delta$) +3.70 +5.19 +3.70 +5.54 +3.62 +23.77 +11.03 +11.27 +4.25
[1pt/2pt] Qwen3-14B 50.94 59.34 50.94 46.92 61.96 91.66 69.88 57.57 12.37
Qwen3-14B (w/ Mate) 53.42 70.21 53.42 47.65 61.40 96.63 72.09 59.99 14.53
Qwen3-14B ($\Delta$) +2.48 +10.87 +2.48 +0.73 -0.56 +4.97 +2.21 +2.42 +2.16
Legend: W. = Weighted; M. = Mean; EMR = Exact Match Ratio; $\Delta$ = Improvement of “(w/ Mate)” over the baseline.
Table 2: Performance comparison on objective evaluation tasks. Mate training significantly improves open-source models’ mistake classification and sentence localization capabilities. The best result is in bold, second-best is underlined.
Experiments & Analysis
The experiments comprise three core components: 1) Experimental Setup: fine-tuning configuration; 2) Model Performance Evaluation: objective evaluation through mistake classification and sentence localization tasks, subjective evaluation via LLM-as-a-Judge and human experts based on professional standards; 3) Downstream Task Validation: verifying model effectiveness through empathy classification and therapist self-efficacy enhancement in real-world mental health applications.
Experimental Setup
We performed parameter-efficient fine-tuning using LoRA (Low-Rank Adaptation) method on NVIDIA H800 GPU. The specific hyperparameter settings were as follows: $r=16$, LoRA scaling factor $\alpha=32$, LoRA dropout rate of $0.05$, the AdamW optimizer, learning rate of 0.0002, batch size of 4, gradient accumulation of 16 and epoch of 2.
Model bleu-4 rouge-1 rouge-2 rouge-l
Llama3.1-8B 18.64 33.63 10.72 22.10
Llama3.1-8B (w/ Mate) 55.73 61.98 42.74 51.15
Qwen3-8B 14.66 35.17 12.49 18.17
Qwen3-8B (w/ Mate) 57.06 63.65 44.84 53.14
Qwen3-14B 13.83 35.31 12.97 16.25
Qwen3-14B (w/ Mate) 54.12 60.88 41.27 49.79
Table 3: BLEU-4 and ROUGE scores for two 8B models and one 14B model before vs. after Mate tuning (Unit: %).
Main Result
Objective Evaluation: Mistake Category Classification & Problematic Sentence Location
Table 2 shows that un-tuned open-source models lag far behind proprietary models in error-type classification, whereas the gap is more manageable in sentence-level localization, which relies more on general language representations. After incorporating Mate data, open-source models achieve remarkable improvements in both tasks. Notably, Qwen3-14B (w/ Mate) surpasses DeepSeek-V3 and closed-source models across multiple metrics (72.09% F1-score vs. Claude-Sonnet-4’s 70.04% on Task 2). This validates the effectiveness of our human-in-the-loop data filter pipeline. This progress stems from Mate’s large-scale, high-quality synthetic samples with explicit feedback signals, which inject domain-specific supervisory knowledge into the models. Nevertheless, both open-source and proprietary models share a “high-recall, low-precision” pattern in sentence localization: they mark as many suspect sentences as possible, capturing most real errors but incurring a high false-positive rate. Future work can mitigate this by imposing confidence thresholds and related strategies to boost precision.
Subjective Evaluation: Supervisory Feedback Generation
With Mate-tuned models achieving 30-40 BLEU and 25+ ROUGE-L improvements (Table 3), we examine whether gains translate to better supervisory content. Following the ”LLM-as-a-Judge” paradigm (Lee et al. 2024b), we use Deepseek-R1-0528 (Guo et al. 2025) for pairwise comparisons on 100 test samples across five criteria: Objectivity & Fairness, Constructiveness & Actionability, Professional Depth, Comprehensiveness, and Clarity & Structure.
As shown in Figure 3, fine-tuning with the Mate dataset significantly enhances critique quality across professional dimensions. The model achieves substantial improvements in core competencies: Objectivity, Professional Depth, and Constructiveness, indicating successful learning of evidence-based, theoretically grounded feedback generation. Comprehensiveness shows moderate improvement (43% win vs. 31% loss), suggesting that grasping broader systemic issues remains challenging.
To validate the effectiveness of the automated evaluation, we compares it with manual evaluation results. Specifically, two psychology experts re-conducted pairwise scoring on the samples. As shown in Figure 4, the results indicate a high overall consistency between the manual evaluation and automated evaluation. These results confirm that our approach effectively transforms the LLM from a generic generator into a specialized supervisory tool.
Figure 3: LLM-as-a-judge evaluation results comparing critiques generated by the Qwen3-8B fine-tuned with Mate dataset (Win) against the base model without fine-tuning (Loss). The chart shows the win, loss, and tie rates across five professional criteria.
Figure 4: High agreement between LLM-as-a-Judge and human evaluation in supervisory feedback assessment.
Feedback Performance on Downstream Tasks
Empathy Classification
To investigate whether training on critical feedback tasks enhances broader psychological competencies, we evaluate models on empathy classification. Both feedback generation and empathy recognition require understanding emotional dynamics and appropriate therapeutic responses, making empathy detection a natural testbed for our therapeutically-trained models. We use the Empathy-Mental-Health (Sharma et al. 2020) dataset (911 samples) to classify whether responses demonstrate empathy in a zero-shot setting.
Table Empathy Classification makes three findings explicit. First, while closed-source GPT-4o delivers the highest F1 (0.447), our Qwen-3-8B model fine-tuned with Mate narrows this gap to less than one point (0.440) and surpasses GPT-4o on both Accuracy and Recall (+0.8 pp). Second, the unfine-tuned base model trails far behind, fine-tuning lifts F1 by 19.0 absolute points (+76%), a gain that is statistically significant. Third, the 8B model appears to strike an optimal balance between learning domain-specific patterns and maintaining generalization capability for downstream transfer.
Looking across open source baselines, Deepseek-R1-0528-685B, which is more than 80x larger, still trails our 8B model on metric. This finding suggests that targeted supervision using psychologically salient critiques can offset the advantage of sheer parameter count in affect sensitive tasks. Taken together, the results show that critical feedback tuning not only works in the original generation setting but also transfers to a zero shot empathy classification scenario, and that Mate endows relatively lightweight models with specialist social emotional competence that even much larger LLMs or their vanilla counterparts lack.
Model Accuracy Precision Recall F1-Score
Closed-Source Models
[1pt/2pt] GPT-4o 42.7 58.5 42.7 44.7
Claude-Sonnet-4 23.6 59.4 23.6 24.8
[1pt/2pt]     Open-Source Models
[1pt/2pt] DeepSeek-R1-0528-685B 31.8 58.1 31.8 34.9
DeepSeek-V3-0324-685B 29.0 55.9 29.0 32.7
Mistral-7B-v0.3 27.9 58.2 27.9 31.1
Qwen2.5-1.5B 14.8 49.6 14.8 8.9
Qwen2.5-7B 25.9 60.7 25.9 26.5
Gemma-3-12B 22.9 63.0 22.9 21.4
Gemma-3-4B 16.4 68.8 16.4 9.9
Qwen-3-8B 24.7 60.6 24.7 25.0
Qwen-3-8B (w/ Mate) 43.5 53.6 43.5 44.0
Qwen-3-14B 22.9 60.5 22.9 21.6
Qwen-3-14B (w/ Mate) 31.6 58.9 31.6 35.2
Table 4: Model Performance on zero-shot Empathy Classification Task. All results are statistically significant ($p<0.001$, Bootstrap test).
Effectiveness of Supervised Feedback on Human
To examine whether automated supervision strengthens novice counsellors’ self-efficacy(Larson and Daniels 1998), we built an online platform where participants conducted 15-minute text-based sessions with an LLM-driven virtual client. After each dialogue, a fine-tuned Supervisor model produced concrete improvement suggestions.
Six counselling-psychology graduate students (3 women, 3 men; average age $M_{\text{age}}=23$; all with fewer than ten hours clinical experience) completed the CASES-R(Hahn et al. 2021) twice: once before reading feedback and again after a five-minute reflection period. We focused on eight items from the exploration–insight and action subscales, each rated from 1 (no confidence) to 6 (complete confidence), yielding 48 paired observations. Scores increased significantly on all eight skills after feedback (10,000 sample bootstrap, 95% CIs). Gains were largest for procedural skills of Direct Guidance and Goal Setting, whereas Focus, already near the ceiling, improved only slightly. Interview data indicated that initial overconfidence in Reflection produced a ceiling effect that dampened its measurable improvement.
Overall, the critical-feedback framework powered by LLMs substantially enhanced novices’ confidence across a broad range of generic and task-specific counselling skills, offering encouraging evidence for the integration of automated supervision into psychotherapist training.
Figure 5: Self-Efficacy of Novice Therapists Before and After Supervised Feedback.
Discussion & Conclusion
This study introduces the innovative LLM-as-a-Supervisor task, aimed at significantly enhancing psychotherapist training efficiency through automated supervision systems. Through deep collaboration with clinical experts, we established comprehensive behavioral guidelines covering 15 categories of typical novice therapist errors and developed a mistake-driven multi-agent dialogue-feedback generation pipeline. This transferable methodology effectively addresses the critical challenge of scarce psychological counseling supervision data. The resulting Mate dataset encompasses three core tasks: problematic sentence localization, mistake classification, and supervisory feedback generation, with data quality ensured through dual safeguards of VGR automated refinement and expert manual review. Experimental results demonstrate that models fine-tuned on the Mate dataset achieve significant improvements in both automated and human evaluations, fully showcasing professional supervisory capabilities. Notably, our 8B-parameter model outperforms multiple closed-source large models on empathy classification tasks, providing strong evidence that targeted supervision training can endow lightweight models with professional emotional understanding capabilities that transcend parameter scale limitations. Self-efficacy assessments of novice counselors reveal that training paradigms incorporating supervisory feedback enhance their confidence levels across multiple counseling skills, providing robust empirical support for integrating automated supervision systems into practical training workflows.